OpenStudy (anonymous):

All the edges of a cube are decreasing at a rate of 6cm per second. a. How fast is the volume changing when each edge is 10cm? b. How is the surface area changing at the same instant?

OpenStudy (anonymous):

ds/dt=-6cm/sec a) V=s^3 dV/dt=3s^2*ds/dt dV/dt=3*100*(-6) dV/dt=-1800cm/sec b) SA=6s^2 dSA/dt=12s*ds/dt dSA/dt=12*10*(-6) dSA/dt=-720cm/sec hope it helped

OpenStudy (anonymous):

Thank you so much! Really helped. :)